OpenStudy (anonymous):

Five marbles are randomly selected with replacement. The probability that a black marble is chosen is 15%. What is the probability that at least 2 marbles are black?

OpenStudy (anonymous):

I don't know what that means. ^^^

OpenStudy (anonymous):

well if one marbel is 15% then two marbels would be 30%

OpenStudy (anonymous):

at least two are black, easier to compute none are black one is black then subtract the result from 1

OpenStudy (anonymous):

That's what I was thinking, but that isn't an option. @skinny23

OpenStudy (anonymous):

that is because it is wrong

OpenStudy (anonymous):

what are the options?

OpenStudy (anonymous):

probability that one is black is 15% probability that one is not black is therefore 84% = .85 probability that none are black out of 5 picks is \(.85^5\)

OpenStudy (anonymous):

probability that exactly one is black is \[5\times .15\times .85^4\]

OpenStudy (anonymous):

your answer is therefore \[1-(.85)^5-5\times (.15)\times (.85)^4\]

OpenStudy (anonymous):

0.16, 0.14, 0.86, 0.84 are my options

OpenStudy (anonymous):

So I would get 0.16? @satellite73